OpenStudy (anonymous):

Between which two integers does the square root of 129 lie? A. 8 and 9 B. 9 and 10 C. 10 and 11 D. 11 and 12

OpenStudy (anonymous):

\[1^2=1,2^2=4,3^2=9,4^2=16,5^2=25,6^2=37\] \[7^2=49,8^2=64,9^2=81,10^2=100,11^2=121,12^2=144\]

OpenStudy (anonymous):

Could you explain that alittle more please @mustafa2014

OpenStudy (anonymous):

Ohh would it be 11 and 12? since 121 and 144 is where 129 is between?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Okay Thank you :)

OpenStudy (anonymous):

np :)